Calculate the speed of a car that travels 75 miles in 3 hours.
A girl rides her bike 15m/s for 20s. How far does she travel?
How long would it take for an object moving at 12m/s to go a distance of 12m?
How much time is required to go 3 miles you can run at the rate of 6.3 miles per hour?

Answer by Alan3354(69443)

Calculate the speed of a car that travels 75 miles in 3 hours.
A girl rides her bike 15m/s for 20s. How far does she travel?
How long would it take for an object moving at 12m/s to go a distance of 12m?
How much time is required to go 3 miles you can run at the rate of 6.3 miles per hour?
-----------------
For all of these:
distance = rate*time
d = r*t
----
r = d/t
To get mi/hr, divide the miles by the hours.
---
t = d/r